# Feature Selection for Functional Data Classification

### 11 Jan 2024 v1

@@ Acknowledgments
APPENDIX
Proofs
Proof of Proposition 1
Note that $h(XB)$ can be expressed as a separable sum, i.e.: $h(XB)=\sum_{i=1}^{n}h(X_{(i)}B)$, with $h(X_{(i)}B)=\log\big{(}1+\exp\big{(}-Y_{i}\cdot(X_{(i)}B)\big{)}\big{)}$, where $Y_{i}\in\{-1,1\}$. Hence, we have (Boyd and Vandenberghe 2004):
@@
Next, from (Boschi et al. 2023), we know:
@@
(ii)   Again, from (Boschi et al. 2023), we know:
@@
(iii)   From (Boschi et al. 2023) we have :
@@
Combing the two equations, we complete the proof.
Algorithms’ hyper-parameters
@@
In all scenarios, FSFC performs a path search on a grid of $100$ different $\lambda_{1}$ values. Specifically, $\lambda_{1}=c_{\lambda}\lambda_{max}$ and $\lambda_{2}=(1-\alpha)\lambda_{1}$, where $\lambda_{max}=0.5\max_{j}\lVert(X^{T}Y)_{(j)}/\omega_{j}\rVert$, ${}c_{\lambda}$ belongs to a grid of $100$ values spaced on a logarithmic scale from $1$ to $0.01$, and $\alpha$ is set to $0.2$. After selecting the model that minimizes a 5-fold cross-validation criterion, the adaptive procedure is implemented as detailed in the main text. For each instance, FSFC starts from an initial value of $\sigma^{0}=0.1c_{\lambda}/\lambda_{max}$ increased by a factor of $\max(\min(5,1+10c_{\lambda}),1.1)$ at each iteration. Choosing $\sigma^{0}$ based on $\lambda_{max}$ and $c_{\lambda}$ is a common practice in the DAL optimization literature (Tomioka, Suzuki, and Sugiyama 2011). The tolerance in Equation (11), employed for determining algorithm convergence, is set to $10^{-4}$, and the parameter $\mu$ in the line-search procedure in Equation (10) is set to $0.2$. Finally, we set $k=5$. Across all examined scenarios and for every feature, five functional principal components are enough to capture more than $95\%$ of the curves’ variability, allowing for a good approximation of the features while maintaining manageable problem dimensions. We explored various values of $\alpha$, $\mu$, and $k$, which yielded comparable outcomes not included in the paper.
Additional figures and tables
@@
Table 2: List of the variables analyzed within the SHARE application. The letter adjacent to the variable name denotes whether it is longitudinal (l), scalar (s), or categorical (c). The letter (a) denotes a scalar variable obtained taking an average across the waves where the values were available. For more detailed information, the reader should consult the SHARE project website: https://share-eric.eu/
@@
Figure 3: On the left, a sample of 50 curves from the first feature of the design matrix $\mathcal{X}$ (top) and the 10 non-zero $\mathcal{B}$ coefficients (bottom) are displayed for the given scenario with $n=300$, $p=800$, $p_{0}=10$. On the right, the SHARE project timeline is depicted, which has been sourced from the SHARE website: https://share-eric.eu/data/data-documentation/waves-overview/

### 05 Mar 2024 v2

@@ Acknowledgements
Appendix A Proofs
A.1 Proof of Proposition 1
Note that $h(XB)$ can be expressed as a separable sum, i.e.: $h(XB)=\sum_{i=1}^{n}h(X_{(i)}B)$, with $h(X_{(i)}B)=\log\big{(}1+\exp\big{(}-Y_{i}\cdot(X_{(i)}B)\big{)}\big{)}$, where $Y_{i}\in\{-1,1\}$. Hence, we have (Boyd & Vandenberghe, 2004):
@@ A.2 Proof of Proposition 2
Next, from (Boschi et al., 2023), we know:
@@
(ii)   Again, from (Boschi et al., 2023), we know:
@@
(iii)   From (Boschi et al., 2023) we have :
@@
Combining the two equations, we complete the proof.
Appendix B Algorithms’ hyper-parameters
@@
In all scenarios, FSFC performs a path search on a grid of $100$ different $\lambda_{1}$ values. Specifically, $\lambda_{1}=c_{\lambda}\lambda_{max}$ and $\lambda_{2}=(1-\alpha)\lambda_{1}$, where $\lambda_{max}=0.5\max_{j}\lVert(X^{T}Y)_{(j)}/\omega_{j}\rVert$, ${}c_{\lambda}$ belongs to a grid of $100$ values spaced on a logarithmic scale from $1$ to $0.01$, and $\alpha$ is set to $0.2$. After selecting the model that minimizes a 5-fold cross-validation criterion, the adaptive procedure is implemented as detailed in the main text. For each instance, FSFC starts from an initial value of $\sigma^{0}=0.1c_{\lambda}/\lambda_{max}$ increased by a factor of $\max(\min(5,1+10c_{\lambda}),1.1)$ at each iteration. Choosing $\sigma^{0}$ based on $\lambda_{max}$ and $c_{\lambda}$ is a common practice in the DAL optimization literature (Tomioka et al., 2011). The tolerance in Equation (11), employed for determining algorithm convergence, is set to $10^{-4}$, and the parameter $\mu$ in the line-search procedure in Equation (10) is set to $0.2$. Finally, we set $k=5$. Across all examined scenarios and for every feature, five functional principal components are enough to capture more than $95\%$ of the curves’ variability, allowing for a good approximation of the features while maintaining manageable problem dimensions. We explored various values of $\alpha$, $\mu$, and $k$, which yielded comparable outcomes not included in the paper.
Appendix C Synthetic data generation
Following other work on functional regression (Parodi et al., 2018), we draw each functional feature $\mathcal{X}_{j}$ and each non-zero regression coefficient curve $\mathcal{B}_{j}$ from a zero-mean Gaussian process with a Matern covariance function (Cressie & Huang, 1999) of the form
$\small\begin{split}C(t,s)=\frac{\eta^{2}}{\Gamma(\nu)2^{\nu-1}}\Bigg{(}\frac{\sqrt{2\nu}}{l}\lvert t-s\rvert\Bigg{)}^{\nu}K_{\nu}\Bigg{(}\frac{\sqrt{2\nu}}{l}\lvert t-s\rvert\Bigg{)}\ ,\end{split}$
where $K_{\nu}$ is a modified Bessel function. We employ point-wise variance $\eta^{2}=1$, range $l=0.25$, and smoothness parameter $\nu=3.5$ for both the $\mathcal{X}$’s and the $\mathcal{B}$’s. Before running FSFC, we standardize each feature individually as $\left(\mathcal{X}_{j}-\text{mean}(t)\right)/\text{sd}(t)$, where $\text{mean}(t)$ and $\text{sd}(t)$ are the point-wise average and standard deviation of all instances computed at $t$. Appendix Figure E1 shows some instances of $\mathcal{X}$ and $\mathcal{B}$ for one specific scenario. The categorical response $Y$ is generated according to the logistic regression procedure: each $Y_{i}$ is drawn from a Bernoulli distribution on $\{-1,1\}$ with probability $\hat{p}_{i}=1/\big{(}1+\exp\big{(}X_{(i)}\hat{B}\big{)}\big{)}$.
Appendix D SHARE experiment settings
In Experiment 2, The cardinalities of the “affected” and “healthy” subject groups, denoted as $n_{a}$ and $n_{h}$, are as follows: $n_{a}=89,102,313,419$ and $n_{h}=1235,1160,747,577$ for diabetes, myocardial infarction, high cholesterol, and hypertension, respectively. To prevent unbalanced scenarios, for each replication, we randomly select $n_{a}$ subjects from the healthy pool, resulting in a total of $n=2n_{a}$ observations evenly balanced between the two classes. The observations are then divided into training and test sets, with $n_{test}=n/4$.
In Experiment 1, each replication uses a training set of only 20 random subjects. All the other observations are part of the testing set.
Appendix E Additional figures and tables
Table E1: Simulation results for the integration of Feature Extraction with LSTM (feLSTM) and SVM (feSVM). For each curve, the following 12 attributes have been extracted: mean, variance, median, maximum, minimum, skewness, kurtosis, zero crossing rate, spectral entropy, spectral kurtosis, area under the curve, and autocorrelation. The table reports the average Classification test accuracy and CPU processing time, measured in seconds, for 50 instances in each scenario.
Accuracy test CPU time   $p_{0}$ feLSTM feSVM feLSTM feSVM   $n=300$ $p=800$ 2 0.535 0.495 67.7 52.6 5 0.532 0.504 67.8 52.9 10 0.538 0.502 67.7 52.6 20 0.539 0.514 67.8 52.6 $n=600$ $p=2000$ 2 0.526 0.504 298.1 271.2 5 0.526 0.509 298.1 273.5 10 0.531 0.504 298.2 271.2 20 0.530 0.505 297.9 271.3
Table E2: SHARE results for the integration of Feature Extraction with LSTM (feLSTM) and SVM (feSVM). For each curve, the extracted attributes are the ones listed in Table E1. The table reports the average classification test accuracy for 100 instances for each response.
feLSTM feSVM   Experiment 1 Diabetes 0.603 0.731 Heart Attack 0.612 0.732 High Colesterol 0.556 0.659 Hypertension 0.614 0.703 Experiment 2 Diabetes 0.652 0.979 Heart Attack 0.632 0.974 High Colesterol 0.596 0.985 Hypertension 0.675 0.987
@@ Appendix E Additional figures and tables
Table E3: List of the variables analyzed within the SHARE application. The letter adjacent to the variable name denotes whether it is longitudinal (l), scalar (s), or categorical (c). The letter (a) denotes a scalar variable obtained taking an average across the waves where the values were available. For more detailed information, the reader should consult the SHARE project website: https://share-eric.eu/
@@
Figure E1: On the left, a sample of 50 curves from the first feature of the design matrix $\mathcal{X}$ (top) and the 10 non-zero $\mathcal{B}$ coefficients (bottom) are displayed for the given scenario with $n=300$, $p=800$, $p_{0}=10$. On the right, the SHARE project timeline is depicted, which has been sourced from the SHARE website: https://share-eric.eu/data/data-documentation/waves-overview/